### v2

# Foundation Model for Composite Materials and Microstructural Analysis

Abstract
The rapid advancement of machine learning has unlocked numerous opportunities for materials science, particularly in accelerating the design and analysis of materials. However, a significant challenge lies in the scarcity and high cost of obtaining high-quality materials datasets. While foundation models pre-trained on large datasets have excelled in fields like natural language processing by leveraging latent features through transfer learning, their application in materials science remains limited. Here, we present a foundation model specifically designed for composite materials. Pre-trained on a dataset of short-fiber composites to learn robust latent features, the model accurately predicts homogenized stiffness during transfer learning, even with limited training data. Additionally, our model effectively predicts the material’s nonlinear behavior by transferring these learned features to an Interaction-based Material Network, which is a constitutive surrogate model. These results demonstrate the potential of our foundation model to capture complex material behaviors. Our findings validate the feasibility and effectiveness of foundation models in composite materials. We anticipate extending this approach to more complex three-dimensional composite materials, polycrystalline materials, and beyond. Moreover, this framework enables high-accuracy predictions even when experimental data are scarce, paving the way for more efficient and cost-effective materials design and analysis.
Keywords Foundation model  $\cdot$ Composite material  $\cdot$ Transfer learning  $\cdot$ Self-supervised learning  $\cdot$ Masked autoencoder
1 Introduction
The mechanical properties of composite materials are highly dependent on their microstructure. Understanding and predicting the effective properties of composites based on their microstructural configurations is crucial for designing and optimizing advanced materials. Machine learning (ML) models have emerged as valuable tools for rapidly predicting these properties, opening new avenues in materials science research [1, 2, 3, 4, 5, 6]. For instance, Abueidda et al. [3] developed a convolutional neural network (CNN) to predict the mechanical properties of two-dimensional checkerboard composites. In addition to CNNs, other architectures have been employed to address different aspects of material property prediction. Mianroodi et al. [4] utilized a U-Net architecture to predict the stress fields in polycrystalline materials, demonstrating the ability of ML models to capture complex stress distributions. Graph neural networks (GNNs) have also been integrated with deep material networks to predict the mechanical responses of composites [5]. Furthermore, Zhongbo et al. [6] employed a pre-trained transformer model to predict the mechanical response of elastoplastic composite materials, illustrating the versatility of transformer architectures in materials modeling.
While these models have achieved considerable success, they predominantly rely on supervised learning, which requires large, labeled datasets for training. The necessity of extensive datasets poses a significant challenge due to the scarcity and high cost of acquiring training data, whether through experiments or simulations. Moreover, many of these models are problem-specific and may not generalize to different tasks, limiting their broader applicability [7]. In contrast, self-supervised learning approaches have revolutionized fields such as natural language processing (NLP) and computer vision by reducing the dependence on labeled data. Notably, models like BERT (Bidirectional Encoder Representations from Transformers) [8, 9] have utilized masked language modeling to learn rich representations from unlabeled text data. The advent of foundation models has further centralized information from diverse data modalities, enabling a single model to be adapted for a wide range of downstream tasks [10, 11].
The concept of foundation models has also been successfully applied to computer vision. Masked autoencoders (MAEs), introduced by He et al. [12], employ a strategy analogous to BERT by masking portions of the input image and training the model to reconstruct the missing parts. In these MAEs, the encoder is replaced with vision transformers (ViTs) [13], and the decoder is a lightweight transformer designed for reconstruction. MAEs have demonstrated performance that, in some instances, outperforms state-of-the-art contrastive learning methods with ViTs, highlighting the efficacy of self-supervised learning in visual domains [14].
Despite these advancements, exploring foundation models within materials science remains limited. Applying self-supervised learning to materials offers the potential to overcome data scarcity and enhance model generalizability across different tasks and material systems. Therefore, in this work, we propose a foundation model specifically designed for composite materials, termed the material masked autoencoder (MMAE), to validate the feasibility of this approach.
To validate the MMAE, we pre-trained the model on a large dataset of composite microstructures and evaluated its performance on two downstream tasks. First, we applied transfer learning to predict the homogenized stiffness components of composite materials, achieving high accuracy even with limited labeled data. Second, we extended the transfer learning framework to predict Interaction-based Material Network (IMN) parameters, enabling the extrapolation of the nonlinear behavior of composite materials [15, 16]. These results demonstrate that the MMAE effectively adapts to various downstream tasks in materials modeling and analysis.
The remainder of the paper is organized as follows: Section 2 outlines the methodology, including data generation and model architecture. Section 3 presents the results and discussion of our experiments. Finally, Section 4 provides concluding remarks and future directions.
2 Methods
2.1 Masked autoencoders
The MMAE employed in this study was primarily based on the MAE architecture proposed by He et al. [12]. As depicted in Figure 1, the MMAE consists of an encoder-decoder structure, where the encoder is a ViT [13] with a patch size of $16\times 16$ pixels, embedding dimension of 256, 12 transformer blocks, and 4 attention heads per block. A linear layer bridges the encoder and decoder, projecting the encoder’s output embeddings to the appropriate dimension for the decoder. The decoder, tailored for reconstruction, comprises 8 transformer blocks with an embedding dimension of 128 and 16 attention heads per block.
Each composite image, originally of size $224\times 224$ grayscale pixels(1 channel), was divided into 196 non-overlapping patches, each measuring $16\times 16$ pixels. These patches were processed using a 2D convolutional layer (Conv2D) to generate token embeddings, which were subsequently sent to the MMAE encoder.
Figure 1: Architecture of the MMAE during pre-training. The input microstructure image is divided into patches, with a subset masked (indicated by striped patterns). The visible patches are embedded and fed into the encoder, and the decoder reconstructs the masked patches based on the encoder’s output.
We applied random masking to a specified proportion of the image patches to introduce the self-supervised learning objective. Specifically, a random subset of patches without replacement was selected to be masked according to a predetermined masking ratio. For example, with a masking ratio of 85%, only 15% (29 out of 196) of the patches remained visible.
The tokens corresponding to the visible patches were fed into the encoder, which processes them through its transformer blocks to produce latent feature representations. Accordingly, the decoder received the encoder’s latent features and positional embeddings for visible and masked patches. It reconstructed the pixel values of the masked patches using its transformer blocks, integrating contextual information from the encoded features. By reconstructing the missing patches, the decoder helps the model learn to infer unseen parts of the microstructure from the available information.
The training objective was to minimize the mean squared error (MSE) between the reconstructed and original pixel values, computed only over the masked patches. By focusing the reconstruction loss solely on the masked patches, the MMAE was encouraged to learn meaningful representations that can accurately predict missing information based on context.
Through this self-supervised approach, the MMAE leverages large amounts of unlabeled data to effectively learn the statistical correlations and spatial patterns inherent in composite microstructures. The encoder learns to generate latent features that encapsulate essential microstructural information, which can be transferred to downstream tasks through fine-tuning or other transfer learning techniques.
2.2 Transfer learning for homogenized stiffness components
Figure 2: Illustration of the MMAE during transfer learning. The encoder processes the input microstructure image, and the [cls] token output is used as a representation. A regression head is connected to the [cls] token to predict the homogenized stiffness components.
To evaluate the effectiveness of the embeddings learned by the pre-trained MMAE for predicting homogenized material properties, we employed transfer learning techniques utilizing the [cls] token output from the encoder as a global representation of the input microstructure. We explored linear probing and fine-tuning approaches, as illustrated in Figure 2.
2.2.1 Linear probing
We performed linear probing to assess the robustness of the latent features extracted by the MMAE. Specifically, we utilized the [cls] token output from the frozen MMAE encoder to represent each input microstructure image. The [cls] token is a unique token added to the sequence of patch embeddings during encoding, designed to capture a global representation of the entire input image. By freezing the encoder weights, we ensured that the learned features remained unchanged during this step.
We trained linear regression models using the [cls] token embeddings as inputs to predict each of the homogenized stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$). The linear regression models were optimized using the MSE loss function. This approach evaluates the effectiveness of the representations learned during pre-training for the downstream task without further modifying the encoder’s parameters.
2.2.2 Fine-tuning
We also used the [cls] token output from the MMAE encoder for fine-tuning as the input representation. However, unlike linear probing, we updated the encoder weights during training. We conducted two types of fine-tuning:
  1. 1.
    End-to-end fine-tuning: In this approach, we allowed all the parameters of the MMAE encoder to be updated during training on the downstream regression tasks. This means the model could adjust its learned representations to better predict homogenized stiffness components. Again, we used the [cls] token to represent the input microstructure and trained a regression head on top of it.
  2. 2.
    Partial fine-tuning: To investigate the impact of fine-tuning only a subset of the model parameters, we performed partial fine-tuning. Specifically, we fine-tuned only the last few transformer blocks of the encoder while keeping the earlier layers frozen. By doing so, we aimed to adjust task-specific representations in the fine-tuned blocks while retaining the general feature representations learned during pre-training in the frozen blocks. This approach allows us to explore a balance between computational efficiency and model performance. The [cls] token was used as the input to the regression head.
The regression head consisted of a fully connected feedforward neural network optimized using the MSE loss function for both fine-tuning methods.
2.3 Transfer learning for IMN parameters
This framework employs transfer learning by fine-tuning the pre-trained MMAE to predict the IMN parameters. The MMAE encodes microstructural information into embeddings, which are passed through a linear projection layer to generate the IMN parameters. During online prediction, the IMN utilizes these parameters to extrapolate nonlinear material behavior, enabling efficient and accurate modeling of complex materials. The schematic of the proposed FM-IMN framework is shown in Figure 3.
Figure 3: Framework of the FM-IMN illustrated for a 2D two-phase RVE. $\mathbf{I}_{s}$ represents the image representation, while $\mathbf{C}^{p1}_{s}$, $\mathbf{C}^{p2}_{s}$, and $\overline{\mathbf{C}}^{\text{DNS}}_{s}$ denote stiffness matrices generated by DNS.
2.3.1 End-to-end offline fine-tuning
The IMN is an efficient physics-based constitutive surrogate model used for multiscale modeling of materials [15, 16, 17]. A notable advantage of the IMN is its ability to be trained using only linear elastic data while achieving accurate nonlinear extrapolation during online prediction. This capability stems from the IMN architecture, which incorporates de-homogenization and homogenization processes during online prediction, enabling it to model complex material behaviors without requiring extensive nonlinear datasets.
The IMN comprises two main components: material nodes and the material network. Specifically, an IMN with a layer depth of $N$ consists of $2^{N}$ material nodes and a perfect binary tree structure called the material network, which contains $12^{N}-$ 1 tree nodes. Each material node represents a material constituent phase and is associated with a trainable parameter $w$ encoding the volume fraction of that phase. Each tree node in the material network has a trainable parameter $\theta$, which represents the stress-equilibrium direction. The collection of all trainable parameters forms $\mathcal{F}$, with a total of $12^{N+1}-$ 1 parameters.
To leverage the microstructural features extracted by the MMAE, we used the [CLS] token output from the encoder, representing a latent embedding of the input microstructure image $I_{s}$. This [CLS] token was passed through a linear projection layer to produce the flattened IMN parameters $\mathcal{F}_{s}$ for each sample $s$. Subsequently, we directly used these parameters $\mathcal{F}_{s}$ to parameterize the IMN, which predicted the homogenized stiffness matrix $\overline{\mathbf{C}}^{\text{MMAE-IMN}}_{s}$, given the constituent stiffness matrices $\mathbf{C}^{p1}_{s}$ and $\mathbf{C}^{p2}_{s}$.
End-to-end fine-tuning was performed to jointly optimize the MMAE encoder and the linear projection layer parameters. To quantify the error between the predicted homogenized stiffness $\overline{\mathbf{C}}^{\text{MMAE-IMN}}_{s}$ and the ground truth $\overline{\mathbf{C}}^{\text{DNS}}_{s}$, we computed the mean squared error (MSE) using the Frobenius norm:
$\text{ERR}=\frac{1}{N_{\text{dataset}}}\sum_{s=1}^{N_{\text{dataset}}}\frac{\left\|\overline{\mathbf{C}}^{\text{DNS}}_{s}-\overline{\mathbf{C}}^{\text{MMAE-IMN}}_{s}(I_{s},\mathbf{C}^{p1}_{s},\mathbf{C}^{p2}_{s})\right\|_{\text{F}}}{\left\|\overline{\mathbf{C}}^{\text{DNS}}_{s}\right\|_{\text{F}}},$ (1)
where $N_{\text{dataset}}$ is the total number of samples, and $\|\cdot\|_{\text{F}}$ denotes the Frobenius norm. We fine-tuned the MMAE encoder and the linear projection layer to accurately predict the homogenized stiffness matrices based on the microstructure images and the constituent material stiffness matrices by minimizing this error.
2.3.2 Online prediction
After training, the MMAE encoder and linear projection layer can infer the IMN parameters $\mathcal{F}$ for any given RVE microstructure image $I$. The inferred IMN, leveraging its online prediction algorithm and the constitutive laws of the constituent phases, performs nonlinear extrapolation to predict the material’s response under diverse loading conditions [15, 16].
By integrating the MMAE’s capability to extract detailed microstructural embeddings with the IMN robust nonlinear extrapolation, this framework provides an efficient and versatile tool for nonlinear material modeling.
2.4 Dataset generation
2.4.1 Generation of microtrsturce
We generated synthetic microstructures for pre-training the MMAE and downstream tasks using the random sequential adsorption (RSA) method [18]. RSA is a stochastic process wherein particles are randomly and sequentially placed into a domain without overlapping, allowing us to create composite microstructures with inclusions of varying quantities, shapes, sizes, and orientations. This method effectively simulates realistic materials systems by capturing the randomness inherent in composite microstructures.
For the short-fiber composites, we considered representative volume element (RVE) containing elliptical inclusions (fibers) dispersed within a matrix. The microstructures were defined within a three-dimensional descriptor space characterized by:
  1. 1.
    Number of Particles ($N_{p}$): The number of fibers in each RVE, ranging from 15 to 35.
  2. 2.
    Aspect Ratio ($A_{r}$): The ratio of the major to minor axes of the elliptical fibers, ranging from 1 to 4.
  3. 3.
    Volume Fraction ($v_{f}$): The fraction of the RVE volume occupied by the fibers, ranging from 10% to 40%.
The actual sizes of the fibers in the RVEs were determined based on $N_{p}$ and $v_{f}$. The orientations of the fibers were assigned randomly during their placement within the RVEs to mimic the stochastic nature of natural materials.
We generated three datasets of RVEs using Latin hypercube sampling [19] for short-fiber composites:
  1. 1.
    Pre-training dataset: A large dataset comprising 100,000 grayscale images of short-fiber composite microstructures was generated for self-supervised pre-training of the MMAE. This dataset captures a wide range of fiber arrangements, aspect ratios, and volume fractions, enabling the MMAE to learn robust and generalizable features from diverse microstructures.
  2. 2.
    Homogenized stiffness dataset: A dataset of 5,000 microstructures was generated for predicting homogenized stiffness components.
  3. 3.
    IMN dataset: A smaller dataset of 300 grayscale images was created for IMN transfer learning to predict nonlinear material behavior.
For the circular-inclusion composites, we generated 5,000 microstructures by uniformly sampling the volume fraction $v_{f}$ within the range of 10% to 40%. The inclusions were circular, and their sizes were fixed. The number of inclusions varied according to the specified volume fractions to maintain the desired $v_{f}$. This dataset was utilized to assess the MMAE’s transfer learning performance on a distinct composite system characterized by different microstructural features. This dataset was used for the downstream task of predicting homogenized stiffness components.
2.4.2 Dataset for homogenized stiffness conponents transfer learning
To obtain the homogenized stiffness components of the generated microstructures, we performed DNS using the RVE module available in LS-DYNA [20]. The simulations were conducted under periodic boundary conditions to accurately capture the effective mechanical response of the composites.
The material properties assigned to the matrix and inclusions were as follows:
  • Matrix: Young’s modulus $E=100$ , Poisson’s ratio $\nu=0.30$.
  • Short-fiber/Circular-inclusion: Young’s modulus $E=500$ , Poisson’s ratio $\nu=0.19$.
Assuming linear elastic material behavior, we focused on predicting homogenized stiffness components: $\bar{C}_{1111}$, $\bar{C}_{2222}$, and $\bar{C}_{1212}$. The DNS computations provided the labels required for training and validation during transfer learning. The dataset was split 80/20 into training and validation subsets.
2.4.3 Dataset for IMN transfer learning
For IMN transfer learning, we used a dataset comprising 300 RVEs of short-fiber composite microstructures. To account for material property variability, we augmented the dataset by sampling 300 unique pairs of stiffness matrices $\mathbf{C}^{p1}_{s}$ and $\mathbf{C}^{p2}_{s}$ for each RVE, representing the matrix (Phase 1) and fiber (Phase 2) materials, respectively. This sampling methodology follows the approach outlined in [5].
For each pair of material properties, DNS was performed using LS-DYNA to compute the corresponding homogenized stiffness matrix $\overline{\mathbf{C}}^{\text{DNS}}_{s}$, resulting in a total of $N_{\text{dataset}}=90{,}000$ instances. The dataset was split into training and validation subsets with an 80/20 ratio, allocating 240 RVEs for training and 60 RVEs for validation.
2.5 Saliency map
We utilized saliency maps to investigate the explainability of MMAE and pinpoint the regions of the microstructure images that have the most significant influence on the model predictions. Saliency maps highlight the input areas that impact the model’s output most, offering visual insights into the model’s decision-making process. They are widely used in various fields, such as image classification and object detection, to assess the relative importance of individual input features in model predictions [21, 22, 23, 24].
In this study, since the task was a regression, the saliency map $M$ was computed based on the gradient of the MSE loss between the predicted value $\hat{y}$ and the ground truth $y$ with respect to the input features. The saliency map for a specific input instance $X_{0}$ is defined as:
$M=\left|\frac{\partial\text{MSE}(y,\hat{y})}{\partial X}\right|_{X_{0}}$ (2)
where $X$ represents the input features, and $X_{0}$ refers to a specific input instance. The absolute value of the gradients $M$ captures the sensitivity of the MSE loss to variations in each input feature, reflecting how changes in the input influence the predicted output.
Larger values in the saliency map $M$ indicate input features that influence the model’s output more. This is because higher gradient magnitudes imply that even small perturbations in these features can lead to significant changes in the MSE loss. By visualizing these saliency maps, we can pinpoint the most influential regions within the microstructure images for predicting the homogenized stiffness components.
3 Results and Discussion
3.1 Pre-trained MMAE reconstruction performance
To assess the pre-trained MMAE’s ability to capture latent features of composite microstructures, we evaluated its reconstruction performance using a high masking ratio of 85%, leaving only 15% (29 out of 196) of image patches visible during training. This challenging setup tests the model’s generalization capabilities on unseen data from downstream task datasets.
As shown in Figure 4, the pre-trained MMAE accurately reconstructs composite microstructure instances, including both short-fiber and circular-inclusion composites, despite the severe masking employed during training. Notably, although the circular-inclusion composites were not included in the pre-training dataset, the MMAE could still reconstruct them effectively. These results demonstrate the MMAE’s robustness in learning representative microstructural features and its capacity to generalize to unseen microstructures, indicating its potential for effective transfer learning in downstream tasks.
Figure 4: Reconstruction of unseen images using the pre-trained MMAE with a masking ratio of 85%. Each triplet includes the ground-truth image (left), the masked image (middle), and the reconstructed image (right). Instances #1 to #4 correspond to circular-inclusion composites, while instances #5 to #8 correspond to short-fiber composites.
3.2 Transfer learning for homogenized stiffness components
To evaluate the effectiveness of the embeddings learned by the pre-trained MMAE, we conducted transfer learning experiments on three downstream regression tasks aimed at predicting the homogenized stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$). We utilized two datasets for these tasks: the short-fiber composite dataset and the circular-inclusion composite dataset.
3.2.1 Linear probing
To assess the utility of the latent features extracted by the MMAE, we performed linear probing by training a linear regression model on top of the fixed pre-trained embeddings to predict the homogenized stiffness components. We investigated the impact of the masking ratio used during pre-training on the linear probing performance for both the short-fiber and circular-inclusion composite datasets.
Figure 5 presents the linear probing performance on the short-fiber composite dataset. The results indicate that the validation set R2 scores decrease significantly as the masking ratio decreases from 35% to 15%, suggesting that very low masking ratios during pre-training may reduce the informativeness of the latent features for linear probing. Nevertheless, the performance remains stable and satisfactory even when the masking ratio exceeds 45%. Notably, the highest validation set R2 score of 0.93 is achieved at a masking ratio of 85%, indicating that higher masking ratios can lead to more effective feature learning during pre-training. These findings highlight the importance of selecting an appropriate masking ratio to ensure the extraction of meaningful representations that generalize well to downstream tasks.
Figure 5: Linear probing performance on the short-fiber composite dataset. (a) The average validation set R2 score of the homogenized stiffness components versus the masking ratio used during pre-training. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the masking ratio.
Figure 6 presents the linear probing results on the circular-inclusion composite dataset. Despite the MMAE being pre-trained on short-fiber composites, the latent features are still effective for predicting the properties of circular-inclusion composites. Similar to the short-fiber composite results, an extremely low masking ratio during pre-training adversely affects performance. However, when the masking ratio exceeds 25%, the performance generally remains high, with average validation set R2 scores exceeding 0.975. The slight dip in performance at a masking ratio of 75% suggests some sensitivity to the masking ratio, but overall, the results demonstrate the generality of the learned embeddings across different composite types.
Figure 6: Linear probing performance on the circular-inclusion composite dataset. (a) The average validation set R2 score of the homogenized stiffness components versus the masking ratio used during pre-training. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the masking ratio.
These findings demonstrate that the choice of masking ratio during pre-training significantly influences the quality of the features extracted by the MMAE for linear probing tasks. A higher masking ratio encourages the model to learn more robust and generalizable representations, which are crucial for accurate predictions without further fine-tuning. This observation aligns with findings in other domains where increased masking ratios in masked autoencoders lead to better feature learning[25, 12, 26].
3.2.2 End-to-end fine-tuning
We employed end-to-end fine-tuning on downstream tasks to evaluate the MMAE’s capacity for modeling nonlinear relationships in the data. Unlike linear probing, which updates only a subset of parameters, end-to-end fine-tuning optimizes the entire parameter set, allowing the model to adapt comprehensively to task-specific challenges.
As shown in Figure 7, the end-to-end fine-tuning performance on the short-fiber composite dataset is relatively insensitive to the masking ratio used during pre-training. In contrast to linear probing, where lower masking ratios led to decreased performance, the end-to-end fine-tuning consistently achieves high average validation set R2 scores of approximately 0.958 across all masking ratios. This result suggests that fine-tuning can effectively adapt the model to the downstream task, regardless of the initial quality of the pre-trained embeddings.
Figure 7: End-to-end fine-tuning performance on the short-fiber composite dataset. (a) The average validation set R2 score of the homogenized stiffness components versus the masking ratio used during pre-training. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the masking ratio.
Similarly, Figure 8 shows the end-to-end fine-tuning performance on the circular-inclusion composite dataset. Consistent with the short-fiber composite dataset results, the performance is relatively unaffected by the masking ratio during pre-training. The model achieves validation set R2 scores exceeding 0.99 across all masking ratios, indicating excellent predictive capabilities after fine-tuning.
Figure 8: End-to-end fine-tuning performance on the circular-inclusion composite dataset. (a) The average validation set R2 score of the homogenized stiffness components versus the masking ratio used during pre-training. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the masking ratio.
These findings suggest that while the initial choice of masking ratio affects the quality of embeddings for linear probing, its impact diminishes during end-to-end fine-tuning. Fine-tuning adjusts the entire model to the specific task, making it less sensitive to the initial masking ratio and enabling better overall performance.
3.2.3 Partial fine-tuning
We conducted partial fine-tuning experiments to explore the MMAE’s behavior in the intermediate regime between linear probing and end-to-end fine-tuning. In partial fine-tuning, only the last few transformer blocks of the model are updated during training on the downstream task, while the earlier blocks remain frozen. This approach allows the model to adjust task-specific representations in the fine-tuned blocks while retaining the general feature representations learned during pre-training in the frozen blocks.
Figure 9 illustrates the performance on the short-fiber composite dataset as a function of the number of blocks fine-tuned. Here, tuning 0 blocks corresponds to linear probing, and tuning all 12 blocks corresponds to end-to-end fine-tuning. The results show that the accuracy improves as more blocks are fine-tuned, indicating better predictive performance. Notably, the performance increases significantly when the last two blocks are fine-tuned and then begins to plateau, suggesting that fine-tuning even a tiny portion of the model’s parameters can capture the necessary task-specific features.
Figure 9: Partial fine-tuning performance on the short-fiber composite dataset with the MMAE pre-trained at a masking ratio of 85%. (a) The average validation set R2 score of the homogenized stiffness components versus the number of blocks fine-tuned. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the number of blocks fine-tuned.
This observation implies that the relationship between the pre-trained embeddings and the homogenized stiffness components exhibits mild nonlinearity. Fine-tuning just a couple of transformer blocks is sufficient to adapt the model to these nonlinear aspects, achieving performance close to full fine-tuning. The substantial improvement in performance with minimal fine-tuning suggests that the earlier layers of the MMAE have learned general features that are transferable across different tasks and datasets. This property is highly desirable in foundation models, as it enables efficient adaptation to new tasks with reduced computational overhead and minimized risk of overfitting.
3.2.4 Model explainability
To explore the explainability of the MMAE during transfer learning, we utilized saliency maps to understand how the model focuses on different regions of the microstructure images when predicting the homogenized stiffness components. Saliency maps highlight the most critical regions of the input that influence the model’s output, providing visual insights into the model’s decision-making process.
We analyzed the saliency maps for both linear probing and end-to-end fine-tuning using the MMAE pre-trained with an 85% masking ratio. Four microstructure images were selected from the validation set of the short-fiber composite dataset for this analysis. For each image, we generated saliency maps corresponding to the predictions of the stiffness components $\bar{C}_{1111}$, $\bar{C}_{2222}$, and $\bar{C}_{1212}$.
Figure 10: Saliency maps for linear probing on the short-fiber composite dataset using the MMAE pre-trained with an 85% masking ratio. Each microstructure image is paired with saliency maps corresponding to the predictions of the stiffness components $\bar{C}_{1111}$, $\bar{C}_{2222}$, and $\bar{C}_{1212}$. Warmer colors indicate regions with higher saliency (greater influence), while cooler ones indicate lower saliency.
As shown in Figure 10, during linear probing, the saliency maps indicate that the model assigns higher importance predominantly to the matrix regions of the microstructures when predicting all stiffness components. This suggests that the model, constrained by linear regression and with frozen encoder weights, relies mainly on the matrix phase to infer the homogenized stiffness components. The emphasis on the matrix implies that the model is not fully capturing the contributions of the fibers and their interactions with the matrix, which are critical to the composite’s mechanical properties. This limitation potentially hinders the model’s predictive accuracy.
Figure 11: Saliency maps for end-to-end fine-tuning on the short-fiber composite dataset using the MMAE pre-trained with an 85% masking ratio. Each microstructure image is paired with saliency maps corresponding to the predictions of the stiffness components $\bar{C}_{1111}$, $\bar{C}_{2222}$, and $\bar{C}_{1212}$. Warmer colors indicate regions with higher saliency, while cooler ones indicate lower saliency.
In contrast, Figure 11 shows that after end-to-end fine-tuning, the saliency maps highlight the matrix-fiber interfaces more prominently. This shift in focus indicates that the fine-tuned model has learned to capture the nonlinear relationships and critical interactions at the interfaces that significantly influence the composite’s overall stiffness. By paying attention to these regions, the model effectively accounts for factors such as fiber orientation and aspect ratio, essential for accurately predicting the homogenized stiffness components.
The change in saliency patterns from emphasizing the matrix regions in linear probing to focusing on the matrix-fiber interfaces after fine-tuning demonstrates how fine-tuning enables the model to adapt its learned representations to capture more complex and relevant microstructural features. This adaptation improves the model’s ability to encode critical aspects of the microstructure, leading to enhanced performance in predicting the homogenized stiffness components compared to linear probing.
3.2.5 Impact of dataset size in transfer learning
To evaluate the influence of dataset size during transfer learning, we selected the MMAE pre-trained with a masking ratio of 85% and varied the number of instances in the fine-tuning dataset. As shown in Figure 12, we observed that the model’s average validation set R2 score is relatively low when the dataset is very small. However, even with a dataset containing only 1000 instances, the model achieves a validation set R2 score above 0.91, demonstrating acceptable performance. As the number of data instances increases, the performance improves, but with diminishing marginal returns. The average R2 score converges when data instances reach approximately 2,500, indicating that adding more data beyond this point yields minimal performance gains.
Figure 12: Impact of dataset size on the performance of transfer learning using the short-fiber composite dataset with the MMAE pre-trained at a masking ratio of 85%. (a) The average validation set R2 score of the homogenized stiffness components versus the number of data instances. (b) Validation set R2 scores for individual stiffness components ($\bar{C}_{1111}$, $\bar{C}_{2222}$, $\bar{C}_{1212}$) versus the number of data instances.
These results highlight the MMAE’s capability to achieve high predictive accuracy even with very limited labeled data, reinforcing its potential for applications where data scarcity is a challenge. This ability to perform well with small datasets is crucial in materials science, where obtaining high-quality experimental data is often costly and time-consuming. The convergence of performance as the dataset size increases further underscores the model’s efficiency in learning the necessary mappings from microstructures to homogenized stiffness components without requiring large datasets.
3.3 Transfer learning for IMN parameters
To explore the effectiveness of the pre-trained MMAE for more complex downstream tasks, we optimized the IMN parameters to predict nonlinear composite behavior. This process involved two key stages: offline training through end-to-end fine-tuning and online prediction. The specific details of these stages are outlined below.
3.3.1 Offline training
In the offline training phase, we employed MMAE models pre-trained with varying masking ratios to investigate their impact on transfer learning performance in the short-fiber composite dataset. Specifically, we performed end-to-end fine-tuning to enhance the latent features for predicting IMN parameters, focusing on how the pre-training masking ratios influenced the fine-tuning process.
As shown in Figure 13, the validation error remained relatively stable, ranging from 6% to 7% across all tested masking ratios for the short-fiber composite dataset. Notably, the lowest validation error occurred at a masking ratio of 50%, which was selected for subsequent online prediction analyses.
Figure 13: Validation error during offline training using MMAE models pre-trained with different masking ratios.
3.3.2 Online prediction
In the online prediction phase, the MMAE–IMN framework was evaluated on four previously unseen RVEs subjected to complex loading conditions. Despite being trained on an offline dataset comprising only 300 RVEs, the pre-trained MMAE demonstrated robust generalization to these new microstructures, showcasing its capability for effective feature extraction. As shown in Figure 14, the predicted stress–strain curves closely align with the DNS results, with the corresponding microstructures depicted as insets. These findings highlight the MMAE–IMN framework’s ability to accurately extract microstructural features and extrapolate nonlinear mechanical behavior, even with a limited fine-tuning dataset.
Figure 14: Predicted stress-strain curves for four unseen RVEs under complex loading conditions. Insets show the corresponding microstructures, illustrating the phase distributions (blue: Phase 1, red: Phase 2).
Table 1 summarizes the material properties of Phase 1 and Phase 2. Phase 1 is modeled with elastoplastic isotropic hardening, whereas Phase 2 is treated purely as an elastic phase.
Table 1: Material properties for Phase 1 and Phase 2.
Young modulus $E$ Poisson ratio $\nu$ Yield stress $\sigma_{y}$ Tangent modulus
Phase 1 100 GPa 0.3 0.1 GPa 5 GPa
Phase 2 500 GPa 0.19
3.4 Summary
The downstream tasks demonstrate that MMAE’s self-supervised pre-training is effective across diverse data-scarce scenarios. For linear probing, the pre-trained embeddings facilitate the extraction of generalizable latent representations of composite microstructures, maintaining robustness even under data-scarce conditions. For more demanding tasks, such as IMN parameter calibration, end-to-end fine-tuning enhances the model’s adaptability, enabling it to meet specific task requirements with high precision.
Saliency analyses reveal that fine-tuning refines the model’s focus on critical fiber–matrix interfaces, improving predictive accuracy for homogenized stiffness components. These findings underscore the utility of self-supervised pre-training as a robust foundation for downstream tasks, while highlighting the importance of tailored adaptations to achieve accurate and efficient material property predictions, particularly in data-limited contexts.
4 Conclusion
This work proposed the MMAE, a foundation model tailored for composite materials, addressing the persistent challenges of data scarcity and transferability in materials science. Through self-supervised pre-training on a large dataset of short-fiber composite microstructures, the MMAE effectively learned robust and representative latent features, providing a solid foundation for diverse downstream applications.
In the first transfer learning scenario, we fine-tuned the pre-trained MMAE to predict the homogenized stiffness of composite materials, achieving an impressive $R^{2}$ score of up to 0.959, even with limited labeled data. This result highlights the model’s effectiveness in overcoming data-scarcity challenges. In the second scenario, the MMAE was fine-tuned to infer IMN parameters, enabling precise prediction of composite material responses and demonstrating its adaptability to tasks requiring the modeling of intricate material behaviors.
These findings highlight the transformative potential of foundation models in materials science, offering scalable and efficient solutions for property prediction and microstructural analysis. Future work will explore extending the MMAE framework to more complex systems, such as three-dimensional composites and polycrystalline materials, as well as incorporating experimental datasets to enhance its applicability in real-world scenarios.
Acknowledgements
This work is supported by the National Science and Technology Council, Taiwan, under Grant 111-2221-E-002-054- MY3 and 112-2221-E-007-028. We are grateful for the computational resources and support from the NTUCE-NCREE Joint Artificial Intelligence Research Center and the National Center of High-performance Computing (NCHC). In addition, we sincerely thank the anonymous reviewers for their constructive comments.
Declarations
The authors declare that there are no competing interests.
Data Availability
The datasets and trained models used in this study are publicly available at the Zenodo repository: https://doi.org/10.5281/zenodo.14062123.
Code Availability
The code used in this study can be accessed at the following GitHub repository: https://github.com/BerryWei/Material_mask_autoencoder.
Author Contributions
Conceptualization and Project Administration: All authors. Investigation and methodology: Ting Ju Wei. Writing—Original Draft: Ting Ju Wei. Review & Editing: All authors. Resources and Funding Acquisition: Chuin-Shan Chen. These author contributions are defined according to the CRediT contributor roles taxonomy.